# GUT-Scale Smooth Hybrid Inflation with a Stabilized Modulus in Light of ACT and SPT Data

### v1

@@ Abstract:
We analyze a generalized framework of smooth F-term hybrid inflation (smFHI) consistent with gauge coupling unification within the Minimal Supersymmetric Standard Model (MSSM). The embedding of the model in two specific Supergravity settings addresses at the same time the $\eta$ problem and the compatibility with the recent ACT or SPT data. The one relies on the choice of a shift-symmetric Kähler potential for the inflaton which revitalizes the SUSY predictions of smFHI, whereas the other employs a Kähler potential associated with an hyperbolic Kähler manifold. An essential role in both our constructions is played by a decoupled superheavy field without superpotential and Kähler potential inspired by string- and D-brane–-based models. Our proposal can be realized for a variety of representations for the Higgs fields involved in smFHI and assures monotonic inflationary potential.
@@ 1 Introduction
A wide stream of works [37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 50, 51, 52, 53, 54, 55, 56, 57, 58, 60, 49, 59, 61, 62, 63, 64, 65, 66, 67, 68, 69, 32, 33, 34, 35, 36] appeared recently trying to reconcile several well-motivated inflationary models with the observational limits above – mainly that of Eq. (3).
@@
In what follows, we present the salient features of our set-up in Sec. 2 and derive the corresponding inflationary potential (Sec. 3). We then outline the observational constraints relevant to our setup (Sec. 4), before proceeding to our numerical analysis (Sec. 5) and concluding remarks (Sec. 6). In Appendices A and B we propose a sample stabilization of the modulus before smFHI and clarify the status of GUT-scale smFHI within mSUGRA.
@@ 2.1 Superpotentials
The field $S$ is a $\mathbb{G}$-singlet left-handed superfield, and $M_{*}$ represents an effective cutoff scale, naturally comparable to the string scale $M_{\rm S}=5\times 10^{17}~{\mbox{\rm GeV}}$. We also explore a simplified version of smFHI [13, 14], where
@@
The superpotential in Eq. (5) is consistent with a continuous $R$ symmetry [19], under which $S$ and $W$ carry the same charge, while $(\bar{\Phi}\Phi)^{p}$ or $\Psi^{q}$ remain neutral. In addition, we impose on $W$ an extra global $\mathbb{Z}_{p}$ or $\mathbb{Z}_{q}$ discrete symmetry, under which $\Phi\rightarrow e^{2\pi i/p}\Phi$ or $\Psi\rightarrow e^{2\pi i/q}\Psi$ for smFHI-I or smFHI-II, respectively. As a result, only the $p$ or $q$ powers of $\bar{\Phi}\Phi$ or $\Psi$ correspondingly are allowed.
@@ 3.1 SUGRA Potential
where the subscript $I~[\bar{J}]$ denotes derivation with respect to (w.r.t) the complex scalar field $X_{I}~[X_{\bar{J}}^{*}]$ which corresponds to the chiral superfield $X_{I}$ with $X_{I}=S,h$ and $\Phi,\bar{\Phi}$ or $\Psi$. The matrix $K^{I\bar{J}}$ is the inverse of the Kähler metric $K_{I\bar{J}}=K_{,X^{I}X^{\bar{J}}}$ – with the symbol $X^{J}$ as subscript denoting derivation with respect to (w.r.t) $X^{J}$ –, i.e., $K^{I\bar{J}}K_{I\bar{L}}=\delta^{\bar{J}}_{\bar{L}}$. Also the F and D terms are
@@
where the $V_{\rm F0}$ is the SUSY part of $V_{\rm F}$ which is given in Sec. 3.2. The remaining part of $V_{\rm F}$ dependents on the SUGRA scenario. Focusing on the $(S,h)$ subspace of our field space we below derive $\langle{K_{I\bar{J}}}\rangle_{\rm I}$ and $\langle{V_{\rm F}-V_{\rm F0}}\rangle_{\rm I}$. Namely,
@@
    $m^{2}_{\bar{\sigma}}=(M^{4}/m_{\rm P}^{2})\left(3+(2\alpha-\beta)^{2}/\beta+(\alpha(3+\alpha)-\beta-1)(\sigma/m_{\rm P})^{2}\right),$ (19)
@@
    $\langle{K_{hh^{*}}}\rangle_{\rm I}=\left(\alpha\sigma^{2}\left(2(\alpha-1)N-{\sigma}^{2}/{m_{\rm P}^{2}}\right)/4N\left(1+\sigma^{2}/2Nm_{\rm P}^{2}\right)^{2}-\beta m_{\rm P}^{2}\right)|\langle{f_{,h}}\rangle_{\rm I}|^{2}$
@@ 3.2 SUSY Contribution
$\langle{\sigma}\rangle=0~~\mbox{and}~~\begin{cases}\langle{\phi}\rangle=2\left(MM_{*}^{(p-1)}\right)^{1/p}&\mbox{for smFHI-I,}\\ \langle{\psi}\rangle=\left(M^{2}M_{*}^{(q-2)}/d_{q}\right)^{1/q}&\mbox{for smFHI-II,}\end{cases}$ (24)
with $d_{q}=|\mbox{\sf Tr}[T_{1}^{q}]|$ serving as a normalization factor that depends on the representation of $\Psi$. At these vacua, the mass of $\sigma$ is found to be
$m_{\rm I}=\left\langle V_{\rm F0,\sigma\sigma}\right\rangle^{1/2}=M^{2}\cdot\begin{cases}\sqrt{2}p\langle{\Phi}\rangle^{-1}&\mbox{for smFHI-I,}\\ q\langle{\Psi}\rangle^{-1}&\mbox{for smFHI-II.}\end{cases}$ (25)
@@
$V_{\rm F0}=M^{4}\begin{cases}\left(1-({\phi}/{\langle{\phi}\rangle})^{2p})^{2}+4p^{2}(\sigma/\langle{\phi}\rangle)^{2}(\phi/\langle{\phi}\rangle)^{2(2p-1)}\right)&\mbox{for smFHI-I,}\\ \left(1-({\psi}/{\langle{\psi}\rangle})^{q})^{2}+q^{2}(\sigma/\langle{\psi}\rangle)^{2}(\psi/\langle{\psi}\rangle)^{2(q-1)}\right)&\mbox{for smFHI-II.}\end{cases}$ (26)
Inspection of $V_{\rm F0}$ shows that it features at least one inflationary valley (with nonzero $\phi$ or $\psi$) smoothly connected to the SUSY vacuum. Already at tree level, these valleys display a built-in slope that drives $\sigma$ towards the vacuum. Consequently, unlike in standard [19] or shifted [20, 21, 22] FHI, radiative corrections are not required to provide the necessary slope, though they may still contribute subdominantly. Let us, finally, note that no problematic domain walls arise from the spontaneous breaking of the global $\mathbb{Z}_{p}$ $\mathbb{Z}_{q}$] symmetry in Eq. (24), since this breaking occurs already during smFHI.
@@
$\langle{V_{\rm F0,\phi}}\rangle_{\rm I}=0~~\Rightarrow~~\langle{\phi}\rangle_{\rm I}=c_{1p}\left(\sigma/\langle{\phi}\rangle\right)^{1/(1-p)}\langle{\phi}\rangle~~\mbox{with}~~c_{1p}=(2p(2p-1))^{(1/2(1-p))}$ (28)
@@ 3.3 Final Form
    which totally eliminates, the quadratic SUGRA contribution to $\eta$ parameter. On the other hand, thanks to the shift symmetry in Eq. (9) all remaining SUGRA corrections in Eq. (34) vanish and therefore the SUSY results on the inflationary observables reveal. The selected $\beta$ value renders the path $\langle{\bar{\sigma}}\rangle_{\rm I}=0$ well stabilized during smFHI since from Eq. (19) with $\alpha=0$ (for simplicity) we obtain
@@ 4.1 Observational Constraints
In the last part of Eq. (39) we assume that smFHI is followed in turn by an oscillatory phase (due to $S$ field) with mean equation-of-state parameter $w_{S\rm rh}$ radiation and matter domination, $T_{\rm rh}$ is the reheat temperature after smFHI. Given that $w_{\rm rh}$ is found by the standard formula $w_{\rm rh}=(n-2)/(n+2)$ for a power-law potential $\varphi^{n}$, we take for our numerics $w_{S\rm rh}=0$ which corresponds precisely to $n=2$, since $V_{\rm F0}$ in Eq. (23) displays a quadratic dependence on $S$. On the other hand, the oscillations of the Higgs fields give rise to a stiff component of the total matter energy-density – with $w_{\rm Hrh}>1/3$ since $n>4$ – which is overshadowed by that of $S$ [84]. Here it is expected that the inflationary energy density is equally distributed between the oscillatory subsystems of $S$ and the Higgs fields. Motivated by implementations [5, 9, 15] of non-thermal leptogenesis, which may follow smFHI, we set $T_{\rm rh}\simeq(10^{8}-10^{9})~{\mbox{\rm GeV}}$. Although not crucial for the resulting magnitude of $N_{\star}$ we mention that we take for the energy-density effective number of degrees of freedom $g_{\rm rh*}=228.75$ inspired by the MSSM spectrum. With these choices, we obtain $N_{\rm I\star}\simeq 48.2-49.5$.
@@ 4.2 Theoretical Considerations
Monotonicity of $V_{\rm I}$. Depending on the values of the coefficients $c_{iK}$ in Eq. (34), the potential $V_{\rm I}$ may either remain monotonic in $\sigma$ or develop a local minimum and maximum. In the scenarios of interest here, $V_{\rm I}$ turns out to be monotonic, avoiding the complications of metastable extrema – see Sec. 5.3.
@@ 5 Results
Recall that the resolution of $\eta$ problem forces us to determine $\beta$ from Eq. (35b) for shSUGRA and $N$ as a function of $\alpha$ and $\beta$ via Eq. (38) for NSUGRA. The unification condition of Eq. (44) in conjunction with Eq. (24) allows us to obtain $M_{*}$ in terms of $M$ as follows
@@
where $d_{q}=2^{-q}15^{-q/2}\left(3\,2^{q}(-M)^{q}+2\,3^{q}M^{q}\right)$ – note that for smFHI-II we confine ourselves in the case with $\mathbb{G}=SU(5)$. Moreover, $M$ can be determined by enforcing Eq. (41) – note that Eq. (39) constrains $\sigma_{\star}$ which is an internal parameter of the inflationary setting. Therefore, the number of the free parameters of our models can be reduced by two.
@@
$\epsilon_{0}=2(2p)^{2}c_{2p}^{2}\left(\frac{m_{\rm P}}{\sigma}\right)^{2}\left(\frac{\sigma}{\langle{\phi}\rangle}\right)^{4p/(1-p)}~~\mbox{and}~~\eta_{0}=\frac{4c_{2p}(1-3p)p}{(p-1)}\left(\frac{m_{\rm P}}{\sigma}\right)^{2}\left(\frac{\sigma}{\langle{\phi}\rangle}\right)^{2p/(1-p)}.$ (46)
@@
$N_{\star}=\frac{(p-1)\sigma_{\star}^{4+2/(p-1)}\langle{\phi}\rangle^{2p/(1-p)}}{8c_{2p}m_{\rm P}^{2}p(2p-1)}~~\Rightarrow~~\sigma_{\star}=2^{\frac{3(p-1)}{4p-2}}\left(\frac{(p-1)\langle{\phi}\rangle^{-2p/(p-1)}}{c_{2p}m_{\rm P}^{2}N_{\star}p(2p-1)}\right)^{\frac{1-p}{4p-2}}.$ (47)
Solving w.r.t $M$ Eq. (41) we find $M$ as follows
@@
where $\sigma_{\star}$ is given by Eq. (47) – its explicit replacement in the expression above yields a rather lengthly final result. Substituting also Eq. (47) into Eq. (42) we find the predictions for $n_{\rm s}$ and $a_{\rm s}$ which become
@@
Table 1: Input and output parameters compatible with Eqs. (39), (41) and (44), within SUSY or shSUGRA for smFHI-I or smFHI-II and selected $p$ or $q$ values respectively.
@@ 5.2 NSUGRA
Here $_{2}F_{1}$ is the Gauss hypergeometric function. Due to these complicate forms of $N_{\star}$ we are not able to solve the expressions above w.r.t $\sigma_{\star}$ and pursue our analytical approach. Assuming, however, that $\sigma_{\star}$ do not deviate a lot from their values in Eq. (47) or (51) for smFHI-I or smFHI-II respectively we can provide analytical approximations that capture the essential behavior of $n_{\rm s}$. Namely we find
@@
Figure 1: Values of $n_{\rm s}$ allowed by Eqs. (39), (41) and (44) versus $\alpha$ for smFHI-I, $\beta=-1$ or 2-2 and $p=2,3$ and 44 – the marginalized joint $68\%$ $95\%$] c.l. regions from P-ACT-LB-BK18 data are depicted by the dark [light] shaded contours (left plot). Resulting values of $(-N_{0})$ versus $M$ for the same $\beta$ and $p$ (right plot). Shown is also the applied color coding in both plots for the various $\beta$ and $p$ values in the legend of the right plot.
This insight can be explicitly verified by carrying out a precise numerical analysis taking into account all the coefficients in Eq. (37a) – (37d). We use as input parameters $\alpha$ and $\beta=-1$ or 2-2 for selected values of $p$ (smFHI-I) or $q$ (smFHI-II). Our results are shown in Fig. 1 for smFHI-I and Fig. 2 for smFHI-II. In the left panel of each figure we present the allowed values of $n_{\rm s}$ versus $\alpha$ whereas in the right panels we display the corresponding values of $(-N_{0})$ as a function of the mass scale $M$. The inner and outer shaded regions represent the marginalized joint 68% and 95% c.l. contours from the P-ACT-LB-BK18 data. The choices of $p$ and $q$ are given in the right plots. For both types of smFHI we observe that there is a plethora of $\alpha$ values which lead to P-ACT-LB-BK18-compatible $n_{\rm s}$ for both chosen $\beta$. Their magnitudes are less than $0.3$ and decrease as $|\beta|$ increases. Also $|N_{0}|$ decreases below unity as $|\beta|$ increases. Therefore, low $p,q$ and $|\beta|$ values are better motivated from the point of view of naturality. The relevant mass scale $M$ turns out to be of order $1~{\mbox{\rm YeV}}$. E.g. Fixing $\beta=-1$ we obtain
@@
Let us emphasize, finally, that the P-ACT-SPT range in Eq. (4) can be obviously accommodated within NSUGRA too. However, the obtained $\alpha$ and $N_{0}$ values are expected to be somehow suppressed and therefore we consider that this framework is more appropriate to fit the P-ACT-LB-BK18 results in Eq. (3). Nonetheless, we can obtain acceptable solutions even for $\alpha=\beta=0$, i.e., without the presence of $\hat{Z}$ and $^{\hat{K}}$ in Eq. (10) at the cost of a very high $N\sim 800$. Although this value is perfectly acceptable within SUGRA, it has no theoretical support and so it can be considered as less natural.

### v2

@@ Abstract:
We analyze a generalized framework of smooth F-term hybrid inflation (smFHI) consistent with gauge coupling unification within the Minimal Supersymmetric Standard Model (MSSM). The embedding of the model in two specific Supergravity settings addresses at the same time the $\eta$ problem and the compatibility with the recent ACT or SPT data. The one relies on the choice of a shift-symmetric Kähler potential for the inflaton which revitalizes the SUSY predictions of smFHI, whereas the other employs a Kähler potential associated with an hyperbolic Kähler manifold. An essential role in both our constructions is played by a decoupled superheavy field without superpotential and Kähler potential inspired by string- and D-brane-based models. Our proposal can be realized for a variety of representations for the Higgs fields involved in smFHI and assures monotonic inflationary potential.
@@ 1 Introduction
A wide stream of works [37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 50, 51, 52, 53, 54, 55, 56, 57, 58, 60, 49, 59, 61, 62, 63, 64, 65, 66, 67, 68, 69, 32, 33, 34, 35, 36, 70, 71, 72, 73, 74, 75] appeared recently trying to reconcile several well-motivated inflationary models with the observational limits above – mainly that of Eq. (3).
@@
In what follows, we present the salient features of our set-up in Sec. 2 and derive the corresponding inflationary potential (Sec. 3). We then outline the observational constraints relevant to our setup (Sec. 4), before proceeding to our numerical analysis (Sec. 5) and concluding remarks (Sec. 6). In Appendices A, B and C we propose a sample stabilization of the modulus before smFHI, we clarify the status of GUT-scale smFHI within mSUGRA and we display a representative realization of reheating.
@@ 2.1 Superpotentials
The field $S$ is a $\mathbb{G}$-singlet left-handed superfield, and $M_{*}$ represents an effective cutoff scale, naturally comparable to the string scale $M_{\rm S}=5\cdot 10^{17}~{\mbox{\rm GeV}}$. We also explore a simplified version of smFHI [13, 14], where
@@
The superpotential in Eq. (5) is consistent with a continuous $R$ symmetry [19], under which $S$ and $W$ carry the same charge, while $(\bar{\Phi}\Phi)^{p}$ or $\Psi^{q}$ remain neutral. In addition, we impose an extra global $\mathbb{Z}_{p}$ or $\mathbb{Z}_{q}$ discrete symmetry, under which $\Phi\rightarrow e^{2\pi i/p}\Phi$ or $\Psi\rightarrow e^{2\pi i/q}\Psi$ for smFHI-I or smFHI-II, respectively. As a result, only the $p$ or $q$ powers of $\bar{\Phi}\Phi$ or $\Psi$ correspondingly are allowed.
@@ 3.1 SUGRA Potential
where the subscript $I~[\bar{J}]$ denotes derivation with respect to (w.r.t) the complex scalar field $X_{I}~[X_{\bar{J}}^{*}]$. Note that we use the same symbol $X_{I}$ for the scalar fields and the corresponding chiral superfield, with $X_{I}=S,h$ and $\Phi,\bar{\Phi}$ or $\Psi$. The matrix $K^{I\bar{J}}$ is the inverse of the Kähler metric $K_{I\bar{J}}=K_{,X^{I}X^{\bar{J}}}$ – with the symbol $X^{J}$ as subscript denoting derivation w.r.t $X^{J}$ –, i.e., $K^{I\bar{J}}K_{I\bar{L}}=\delta^{\bar{J}}_{\bar{L}}$. Also the F and D terms are
@@
where $V_{\rm F0}$ is the SUSY part of $V_{\rm F}$ which is given in Sec. 3.2. The remaining part of $V_{\rm F}$ dependents on the SUGRA scenario. Focusing on the $(S,h)$ subspace of our field space we below derive $\langle{K_{I\bar{J}}}\rangle_{\rm I}$ and $\langle{V_{\rm F}-V_{\rm F0}}\rangle_{\rm I}$. Namely,
@@
    $m^{2}_{\bar{\sigma}}=(M^{4}/m_{\rm P}^{2})\left(3-(2\alpha-\beta)^{2}/\beta+((3+\alpha)\alpha-\beta-1)(\sigma/m_{\rm P})^{2}\right),$ (19)
@@
    $\langle{K_{hh^{*}}}\rangle_{\rm I}=\left(\alpha\sigma^{2}\left(2(\alpha-1)N-{\sigma}^{2}/{m_{\rm P}^{2}}\right)/4N^{2}\left(1+\sigma^{2}/2Nm_{\rm P}^{2}\right)^{2}-\beta m_{\rm P}^{2}\right)|\langle{f_{,h}}\rangle_{\rm I}|^{2}$
@@ 3.2 SUSY Contribution
$\langle{\sigma}\rangle=0~~\mbox{and}~~\begin{cases}\langle{\phi}\rangle=2\left(MM_{*}^{p-1}\right)^{1/p}&\mbox{for smFHI-I,}\\ \langle{\psi}\rangle=\left(M^{2}M_{*}^{q-2}/d_{q}\right)^{1/q}&\mbox{for smFHI-II,}\end{cases}$ (24)
with $d_{q}=|\mbox{\sf Tr}[T_{1}^{q}]|$ serving as a normalization factor that depends on the representation of $\Psi$. At these vacua, the masses of $\sigma$ and $\phi$ or $\psi$ are found to be equal with common value
$m_{\rm I}=M^{2}\cdot\begin{cases}\sqrt{2}p\langle{\Phi}\rangle^{-1}&\mbox{for smFHI-I,}\\ q\langle{\Psi}\rangle^{-1}&\mbox{for smFHI-II.}\end{cases}$ (25)
@@
$V_{\rm F0}=M^{4}\begin{cases}\left((1-({\phi}/{\langle{\phi}\rangle})^{2p})^{2}+4p^{2}(\sigma/\langle{\phi}\rangle)^{2}(\phi/\langle{\phi}\rangle)^{2(2p-1)}\right)&\mbox{for smFHI-I,}\\ \left((1-({\psi}/{\langle{\psi}\rangle})^{q})^{2}+q^{2}(\sigma/\langle{\psi}\rangle)^{2}(\psi/\langle{\psi}\rangle)^{2(q-1)}\right)&\mbox{for smFHI-II.}\end{cases}$ (26)
Inspection of $V_{\rm F0}$ shows that it features at least one inflationary valley (with nonzero $\phi$ or $\psi$) smoothly connected to the SUSY vacuum. Already at tree level, these valleys display a built-in slope that drives $\sigma$ towards the vacuum. Consequently, unlike standard [19] or shifted [20, 21, 22] FHI, radiative corrections are not required to provide the necessary slope, though they may still contribute subdominantly. Let us, finally, note that no problematic domain walls arise from the spontaneous breaking of the global $\mathbb{Z}_{p}$ or $\mathbb{Z}_{q}$ symmetry in Eq. (24), since this breaking occurs already during smFHI.
@@
$\langle{V_{\rm F0,\phi}}\rangle_{\rm I}=0~~\Rightarrow~~\langle{\phi}\rangle_{\rm I}=c_{1p}\left(\sigma/\langle{\phi}\rangle\right)^{1/(1-p)}\langle{\phi}\rangle~~\mbox{with}~~c_{1p}=(2p(2p-1))^{(1/2(1-p))},$ (28)
@@ 3.3 Final Form
    which totally eliminates, the quadratic SUGRA contribution to $\eta$ parameter. On the other hand, thanks to the shift symmetry in Eq. (9) all remaining SUGRA corrections in Eq. (34) vanish and, therefore, the SUSY results on the inflationary observables reveal. The selected $\beta$ value renders the path $\langle{\bar{\sigma}}\rangle_{\rm I}=0$ well stabilized during smFHI since from Eq. (19) with $\alpha=0$ (for simplicity) we obtain
@@ 4.1 Observational Constraints
In the last part of Eq. (39) we assume that smFHI is followed in turn by an oscillatory phase (due to $S$) with mean equation-of-state parameter $w_{\rm rh}$, radiation and matter domination. Also $T_{\rm rh}$ is the reheat temperature after smFHI. Given that $w$ is found by the standard formula [1]
$w_{\varphi}=(n-2)/(n+2)~~\mbox{for a power-law potential}~~\varphi^{n},$ (41)
we take for our numerics $w_{\rm rh}=w_{S}=0$ which corresponds precisely to $n=2$, since $V_{\rm F0}$ in Eq. (23) displays a quadratic dependence on $S$ for $\phi$ close to its v.e.v in Eq. (24). On the other hand, the oscillations of the Higgs fields give rise to a stiff component of the total matter energy-density – with $w_{\rm H}>1/3$ since $n>4$ – which is overshadowed by that of $S$ [90]. This intuitive picture is verified and further explained through a numerical example presented in Appendix C. Motivated by implementations [5, 9, 15] of non-thermal leptogenesis [91], which may follow smFHI, we set $T_{\rm rh}\simeq(10^{8}-10^{9})~{\mbox{\rm GeV}}$. Although not crucial for the resulting magnitude of $N_{\star}$, we mention that we take for the energy-density effective number of degrees of freedom $g_{\rm rh*}=228.75$ inspired by the MSSM spectrum. With these choices, we obtain $N_{\rm I\star}\simeq 48.2-49.5$.
@@ 4.2 Theoretical Considerations
Monotonicity of $V_{\rm I}$. Depending on the values of the coefficients $c_{iK}$ in Eq. (34), the potential $V_{\rm I}$ may either remain a monotonic function of $\sigma$ or develop a local minimum and maximum. In the scenarios of interest here, $V_{\rm I}$ turns out to be monotonic, avoiding the complications of metastable extrema – see Sec. 5.3.
@@ 5 Results
Recall that the resolution of the $\eta$ problem forces us to determine $\beta$ from Eq. (35b) for shSUGRA and $N$ as a function of $\alpha$ and $\beta$ via Eq. (38) for NSUGRA. The unification condition of Eq. (45) by virtue of Eq. (24) allows us to obtain $M_{*}$ in terms of $M$ as follows
@@
where $d_{q}=2^{-q}15^{-q/2}\left(3\cdot 2^{q}(-M)^{q}+2\cdot 3^{q}M^{q}\right)$ – note that for smFHI-II we confine ourselves in the case with $\mathbb{G}=SU(5)$. Moreover, $M$ can be determined by enforcing Eq. (42) – note that Eq. (39) constrains $\sigma_{\star}$ which is an internal parameter of the inflationary setting. Therefore, the number of the free parameters of our models can be reduced by two.
@@
$\epsilon_{0}=2p^{2}c_{2p}^{2}\left(\frac{m_{\rm P}}{\sigma}\right)^{2}\left(\frac{\sigma}{\langle{\phi}\rangle}\right)^{4p/(1-p)}~~\mbox{and}~~\eta_{0}=\frac{4c_{2p}(1-3p)p}{(p-1)}\left(\frac{m_{\rm P}}{\sigma}\right)^{2}\left(\frac{\sigma}{\langle{\phi}\rangle}\right)^{2p/(1-p)}.$ (47)
@@
$N_{\star}=\frac{(p-1)\sigma_{\star}^{4+2/(p-1)}\langle{\phi}\rangle^{2p/(1-p)}}{4c_{2p}m_{\rm P}^{2}p(2p-1)}~~\Rightarrow~~\sigma_{\star}=2^{\frac{(p-1)}{2p-1}}\left(\frac{(p-1)\langle{\phi}\rangle^{-2p/(p-1)}}{c_{2p}m_{\rm P}^{2}N_{\star}p(2p-1)}\right)^{\frac{1-p}{4p-2}}.$ (48)
Solving Eq. (42) w.r.t $M$ we find
@@
where $\sigma_{\star}$ is given by Eq. (48) – its explicit replacement in the expression above yields a rather lengthly final result. Substituting also Eq. (48) into Eq. (43) we arrive at the predictions for $n_{\rm s}$ and $a_{\rm s}$ which become
@@
Table 1: Input and output parameters compatible with Eqs. (39), (42) and (45), within shSUGRA (or SUSY) for smFHI-I or smFHI-II and selected $p$ or $q$ values respectively.
@@ 5.2 NSUGRA
Here $_{2}F_{1}$ is the Gauss hypergeometric function. Due to these complicate forms of $N_{\star}$ we are not able to solve the expressions above w.r.t $\sigma_{\star}$ and pursue our analytical approach as in Sec. 5.1. Assuming, however, that the $\sigma_{\star}$ values do not deviate a lot from their values in Eq. (48) or (52) for smFHI-I or smFHI-II respectively we can provide analytical approximations that capture the essential behavior of $n_{\rm s}$. Namely we find
@@
Figure 1: Values of $n_{\rm s}$ allowed by Eqs. (39), (42) and (45) versus $\alpha$ for smFHI-I, $\beta=-1$ or 2-2 and $p=2,3$ and $4$ – the marginalized joint $68\%$ $95\%$] c.l. regions from P-ACT-LB-BK18 data are depicted by the dark [light] shaded contours (left plot). Resulting values of $(-N_{0})$ versus $M$ for the same $\beta$ and $p$ (right plot). Shown is also the applied color coding in both plots for the various $\beta$ and $p$ values in the legend of the right plot.
This insight can be explicitly verified by carrying out a precise numerical analysis taking into account all the coefficients in Eq. (37a) – (37d). We use as input parameters $\alpha$ and $\beta=-1$ or 2-2 for selected values of $p$ (smFHI-I) or $q$ (smFHI-II). Our results are shown in Fig. 1 for smFHI-I and Fig. 2 for smFHI-II. In the left panel of each figure we present the allowed values of $n_{\rm s}$ versus $\alpha$ whereas in the right panels we display the corresponding values of $(-N_{0})$ as a function of the mass scale $M$. The inner and outer shaded regions represent the marginalized joint 68% and 95% c.l. contours from the P-ACT-LB-BK18 data. The choices of $p$ and $q$ are given in the right plots. For both types of smFHI we observe that there is a plethora of $\alpha$ values which lead to P-ACT-LB-BK18-compatible $n_{\rm s}$ for both chosen $\beta$. Their magnitudes are less than $0.3$ and decrease as $|\beta|$ increases. Also $|N_{0}|$ decreases below unity as $|\beta|$ increases. Therefore, low $p,q$ and $|\beta|$ values are better motivated from the point of view of naturality. The relevant mass scale $M$ turns out to be of order $1~{\mbox{\rm YeV}}$. E.g., fixing $\beta=-1$ we obtain
@@
Let us emphasize, finally, that the P-ACT-SPT range in Eq. (4) can be obviously accommodated within NSUGRA too. However, the obtained $\alpha$ and $N_{0}$ values are expected to be somehow suppressed and therefore we consider that this framework is more appropriate to fit the P-ACT-LB-BK18 results in Eq. (3). Also, we can obtain acceptable results even for $\alpha=\beta=0$, i.e., without the presence of $\hat{Z}$ and $^{\hat{K}}$ in Eq. (10) at the cost of a very high (positive) $N\sim 800$. Although this value is perfectly acceptable within SUGRA, it has no theoretical motivation and so it can be considered as less natural.